OpenStudy (anonymous):

help Simplify. (2+1/a)/(2/a-a)

OpenStudy (imstuck):

Does that look like this:\[\frac{ 2+\frac{ 1 }{ a } }{ \frac{ 2 }{ a }-a }\]

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

ok so you need to do the numerator first and get the common denominator of a throughout the top, then do the same with the bottom part of the fraction...get a common denominator, which is the same as the top, then combine the whole fraction and get the answer for a. I know that sounds really confusing, so let's start with the top of that mess, the 2 + 1/a.

OpenStudy (imstuck):

Getting a common denominator gives you 2a/a + 1/a which comes to 2a+1/a. And on the bottom you need to common denom of a as well, so that gives you 2/a - a^2/a, which combines to 2-a^2/a. Now you have this:

OpenStudy (imstuck):